OpenStudy (anonymous):

1-(cos^2x)/(1+sinx) write as a single trig function

OpenStudy (dumbcow):

\[\frac{1-\cos^2 x}{1+\sin x} = \frac{\sin^2 x}{1+\sin x}\] multiply by conjugate \[\frac{\sin^2 x}{1+\sin x}*\frac{1-\sin x}{1-\sin x} = \frac{\sin^2 x (1-\sin x)}{1-\sin^2 x} = \frac{\sin^2 x(1-\sin x)}{\cos^2 x}\] \[= \tan^2 x (1-\sin x)\] hmm thats as far as i can go i think